# Attributing the point symmetric structure of core-collapse supernova remnant N132D to the jittering jets explosion mechanism

### v1

@@ 3.2 Previously identified point-symmetric morphology
In Figure 2 I present two panels from Morse et al. (1995) of SNR N132D. Panel a presents in colors the velocity map by [O iii]5007 of fast oxygen-rich filaments and in gray the low-velocity oxygen emission. Panel b presents the O-rich filaments with the four lines that Morse et al. (1995) added to the image by connecting regions that show a symmetric distribution about a common center. They identified a point-symmetric morphological region in the center (although they do not use this term). They identified the intersection of these four lines as the center of the inner, fast, oxygen-rich filament distribution, which they marked by "" in panel a. The "" symbol in panel a marks the center of the outer diffuse oxygen emission, which they determined by fitting an ellipse to that gas. These two centers also appear in panel c of Figure 1.
I added to panel b of Figure 2 the two double-sided arrows from Figure 1, that are the short-symmetry axis that I identify, and the perpendicular axis, which is along the long dimension of SNR N132D. From Figures 1 and 2 I notice the following. (1) Although the two axes I identify do not coincide with those of Morse et al. (1995), the perpendicular axis is also along the long dimension of the fast oxygen filaments in the center. (2) The center I identify by the two double-sided arrows (orange-dot on panels c and d of Figure 1) is close to the center of the inner oxygen filaments that Morse et al. (1995) identify.
@@ 3.3 The relation to the oxygen 3D structure
Vogt & Dopita (2011) map the [O III]5007 dynamics of SNR N132D and reconstruct its 3D structure. They found that the majority of the ejecta form a ring (torus-like) of in diameter. Lasker (1980) already identified the ring structure of the oxygen-rich ejecta near the center. Vogt & Dopita (2011) speculated that the oxygen-rich ring is in the equatorial plane of a bipolar explosion and that the morphology is also strongly influenced by the CSM. In panel a of Figure 3, I present an image adapted from Rho et al. (2023) that shows the Doppler shift measurements of Morse et al. (1995) on top of the general image of SNR N132D in the IR and X-ray. I added the three axes from panel d of Figure 1. The IR emission (diffuse blue regions) reveals material surrounding SNR N132D; however, I do not attribute the structure of the inner oxygen-rich ring (torus) to the CSM or ISM that are located outside the main SNR shell.
@@
Additionally, Vogt & Dopita (2011) associated the fast, oxygen-rich runaway knot (RK; panel b of Figure 2) with a polar jet. I accept this interpretation that polar jets shaped the ring (torus) of oxygen-rich ejecta in the inner part of SNR N132D. I suggest that a pair of opposite jets along the short-symmetry axis participated in the explosion process of SNR N132D; there were several more pairs of jets according to the JJEM. In panel b of Figure 3, I present the three axes I identified in Figure 1 on top of an image adapted from Law et al. (2020). The comparison in the panels of Figure 3 strengthens the relation I found from Figure 2 between the large-scale symmetry of SNR N132D and the symmetry of the oxygen-rich ejecta in the inner region.
@@
The 3D structure reconstruction by Law et al. (2020) is broadly consistent with the 3D geometry constructed by Vogt & Dopita (2011). Law et al. (2020) find the majority of the bright oxygen ejecta to reside in a broken and distorted torus tilted to the plane of the sky and with a radius of , where is the distance to SNR N132D. The velocity along the line of sight is from to . They find the Doppler velocity of the runaway knot to be and its total space velocity , about twice the bulk velocity of the oxygen-rich ejecta of . Their finding supports the argument of Vogt & Dopita (2011) for a polar jet. Of high significance to my claim for the JJEM is Law et al. (2020) finding that the runaway knot is nearly perpendicular to the torus plane and coincident with an X-ray emission spot (inset of panel b of Figure 3) that is substantially enhanced in silicon and somewhat in sulfur relative to the Large Magellanic Cloud and N132D’s bulk ejecta. The different composition from the rest of the ejecta shows that the runaway knot is not a result of an instability, e.g., a finger of Rayleigh-Taylor instability that keeps the same composition stratification of the rest of the ejecta; the runaway knot must be a separate ejection event, namely, a jet. This is the same argument on why the silicon-rich jet of Cassiopeia A cannot be an instability, and it is a jet launched during the explosion process (e.g., Soker 2017).
@@
Law et al. (2020) find some departures from pure axi-symmetry of the torus in that the redshifted material is elevated above the midplane of the torus, while the blueshifted side is below the midplane. In the JJEM, this is part of the point-symmetric morphology that pairs of jets shape during the explosion process. Law et al. (2020) also find that the runaway knot is to the normal to the torus.
@@
Law et al. (2020) identified a break in the torus; I present their identification in Figure 4. Examining some images from their analysis, I identify a counter break, i.e., a break opposite to the one they identified, as shown in panel a of Figure 4; it has a narrower opening. In the same panel, I draw a line from the break through the center and to the counter break that I identify here. As far as I can tell, this symmetry line through the breaks in the torus coincides on the plane of the sky with the perpendicular axis that I draw in Figures 1 and 2, up to the uncertainty in the exact location of the center of explosion. I suggest that a jet, one of two jets in a pair, shaped the break in the torus. The continuation of material outward on the sides of the break supports such an interpretation. An opposite jet shaped the counter break. On panel b of Figure 4, I added the perpendicular axis, but shifted to the center that Law et al. (2020) uses. This axis is aligned along the extension of the oxygen-rich material to the northeast, highlighting the relationship between the large-scale symmetry and the oxygen-rich material.
@@
Law et al. (2020) compare the torus they reconstruct in SNR N132D with that of Cassiopeia A. I note that the torus of Cassiopeia A possesses a rich point-symmetric morphology (Bear & Soker, 2025). I here argue that the torus of SNR N132D also possesses a point-symmetric morphology. Law et al. (2020) compared the runaway knot of SNR N132D to fast ejecta clumps in the Vela SNR, which is another CCSNR with a prominent and rich point-symmetric morphology that only the JJEM can account for (Soker 2023; Soker & Shishkin 2025).
@@ 4 Discussion and Summary
By visually inspecting images of CCSNR N132D, a method I justified in Section 2, I identified a point-symmetric morphology in CCSNR N132D, composed of two symmetry axes. The first one is the short symmetry axis, extending from the northwest ear and through the center of the iron-rich emission that Foster et al. (2025) presented in a recent study (panel a of Figure 1). The second, the perpendicular axis on the plane of the sky, is an axis perpendicular to the short symmetry axis and passing through its center, which is also aligned along the long dimension of SNR N132D; it coincides with the extension of the oxygen-rich gas to the northeast (panel d of Figure 1, Figure 2 and 3, and panel b of Figure 4). The perpendicular axis seems to coincide with the line from the break in the torus that Law et al. (2020) identify to the counter break that I identify (panel a of Figure 4). There might be a third axis: Law et al. (2020) find the redshifted material in the torus to be somewhat above the midplane of the torus and the blueshifted side to be below the midplane. A line between these two regions, that is inclined to the plane of the torus, might define a third symmetry axis in SNR N132D. This deserves further study.
SNR N132D is surrounded by material (CSM and/or ISM), as indicated by the diffuse blue zones in panel a of Figure 3. SNR N132D interacts with the surrounding clouds (e.g., Dopita et al. 2018; Sano et al. 2020; H. E. S. S. Collaboration et al. 2021; Gu et al. 2025). This interaction affects the morphology of the outer regions, but cannot shape the inner oxygen-rich gas. The fact that the inner and outer morphologies correlate shows that the point symmetry is a result of the explosion.
@@
I considered the possibility that the two opposite jets along the long dimension of SNR N132D were not exactly opposite, i.e., a bent pair of jets. I mark the two suggested jets’ directions with solid-orange arrows in panel d of Figure 1; they are at to each other. Shishkin et al. (2025) suggested that one of the pairs of jets that exploded and shaped CCSNR S147 was a bent pair.
@@
The JJEM accounts for the point-symmetric morphology of SNR N132D, as well as for the perpendicular two axes and the bent pair of jets. The perpendicular axes might result from the following two effects. In an early energetic pair of jets, one jet carries much more momentum (and energy) than the other. As a result of that, the NS acquires a kick velocity; this is the kick-BEAP (kick by early asymmetrical pair) mechanism (Bear et al., 2024; Shishkin et al., 2025). The kick velocity is along the axis of the jets. The kick velocity imparts an angular momentum component to the material that the NS accretes later from the collapsing core that is perpendicular to the kick velocity. If this component dominates, then the pair of jets that the accreted gas launches along its angular momentum axis tends to be perpendicular to the kick velocity (e.g., Bear & Soker 2018), i.e., perpendicular to the axis of the first pair of jets. This effect of post-kick accretion might explain the avoidance of small angles between the NS kick velocity and the main jet axis in some CCSNRs (e.g., Soker 2022b; Bear & Soker 2023). Many of the accretion episodes through the intermittent accretion disks in the JJEM last for a time scale that is shorter, or not much longer than the relaxation time of the accretion disk (Soker, 2024c). Because accretion of material with stochastic angular momentum and density fluctuations forms the accretion disk, the two sides of the unrelaxed accretion disk might be unequal in size and structure. Earlier studies of the JJEM followed (Soker, 2024c) and considered that opposite jets differ in their power and opening angle. I here add to the claim by Shishkin et al. (2025) that, in addition, the two jets might not be exactly opposite to each other, and the angle between them can be , e.g., in SNR S147.
@@
SNR N132D is the fifteenth CCSNR with an identified point-symmetric morphology attributed to the JJEM (Section 2). Presently, the point-symmetric morphology is the only property of CCSNe and CCSNRs that can clearly distinguish between the JJEM and the delayed neutrino explosion mechanism (Soker, 2024b). Other properties, such as light curves and nucleosynthesis, are similar but not identical between the two explosion mechanisms. However, the differences in these properties are not at a level that allows for a clear observational determination of the explosion mechanism (Soker, 2024b). The point-symmetrical morphology due to point-symmetric explosion is a robust prediction of the JJEM (although it will not be observed in all CCSNRs) that has no explanation in the neutrino-driven mechanism (e.g., Soker & Shishkin 2025). Some supporters of the neutrino-driven explosion mechanism have noted that their simulations do not produce jets (I attributed it to the limited ability of existing CCSN hydrodynamical numerical codes to handle magnetic field reconnection; Soker 2024a). As a result, they often do not engage with or cite literature related to the JJEM model. This approach, where theoretical expectations (but only by some researchers) are used to dismiss observational evidence, contrasts with the standard scientific methodology in which models are tested and refined based on empirical data. In contrast, the JJEM model argues that the observed point symmetric morphologies in CCSNRs provide significant evidence that favors jet-driven processes over neutrino-driven explosions as the primary explosion mechanism. I therefore encourage the CCSNe research community to consider evaluating the JJEM on equal ground with the delayed neutrino explosion mechanism.
@@ Acknowledgements
I thank Ealeal Bear, Dima Shishkin, and Amit Kashi for their many discussions and comments, which were relevant to the topic of this study. I thank the Charles Wolfson Academic Chair at the Technion for the support.

### v2

@@ 3.2 Previously identified point-symmetric morphology
In Figure 2 I present two panels from Morseetal1995 of SNR N132D. Panel (a) presents in colors the velocity map by [O iii]5007 of fast oxygen-rich filaments and in gray the low-velocity oxygen emission. Panel (b) presents the O-rich filaments with the four lines that Morseetal1995 added to the image by connecting regions that show a symmetric distribution about a common center. They identified a point-symmetric morphological region in the center (although they do not use this term). They identified the intersection of these four lines as the center of the inner, fast, oxygen-rich filament distribution, which they marked by "" in panel (a). The "" symbol in panel (a) marks the center of the outer diffuse oxygen emission, which they determined by fitting an ellipse to that gas. These two centers also appear in panel (c) of Figure 1.
I added to panel (b) of Figure 2 the two double-sided arrows from Figure 1, that are the short-symmetry axis that I identify, and the perpendicular axis, which is along the long dimension of SNR N132D. From Figures 1 and 2 I notice the following. (1) Although the two axes I identify do not coincide with those of Morseetal1995, the perpendicular axis is also along the long dimension of the fast oxygen filaments in the center. (2) The center I identify by the two double-sided arrows (orange-dot on panels c and d of Figure 1) is close to the center of the inner oxygen filaments that Morseetal1995 identify.
@@ 3.3 The relation to the oxygen 3D structure
VogtDopita2011 map the [O III]5007 dynamics of SNR N132D and reconstruct its 3D structure. They found that the majority of the ejecta form a ring (torus-like) of in diameter. Lasker1980 already identified the ring structure of the oxygen-rich ejecta near the center. VogtDopita2011 speculated that the oxygen-rich ring is in the equatorial plane of a bipolar explosion and that the morphology is also strongly influenced by the CSM. In panel (a) of Figure 3, I present an image adapted from Rhoetal2023 that shows the Doppler shift measurements of Morseetal1995 on top of the general image of SNR N132D in the IR and X-ray. I added the three axes from panel (d) of Figure 1. The IR emission (diffuse blue regions) reveals material surrounding SNR N132D; however, I do not attribute the structure of the inner oxygen-rich ring (torus) to the CSM or ISM that are located outside the main SNR shell.
@@
Additionally, VogtDopita2011 associated the fast, oxygen-rich runaway knot (RK; panel b of Figure 2) with a polar jet. I accept this interpretation that polar jets shaped the ring (torus) of oxygen-rich ejecta in the inner part of SNR N132D. I suggest that a pair of opposite jets along the short-symmetry axis participated in the explosion process of SNR N132D; there were several more pairs of jets according to the JJEM. In panel (b) of Figure 3, I present the three axes I identified in Figure 1 on top of an image adapted from Lawetal2020. The comparison in the panels of Figure 3 strengthens the relation I found from Figure 2 between the large-scale symmetry of SNR N132D and the symmetry of the oxygen-rich ejecta in the inner region.
@@
The 3D structure reconstruction by Lawetal2020 is broadly consistent with the 3D geometry constructed by VogtDopita2011. Lawetal2020 find the majority of the bright oxygen ejecta to reside in a broken and distorted torus tilted to the plane of the sky and with a radius of , where is the distance to SNR N132D. The velocity along the line of sight is from to . They find the Doppler velocity of the runaway knot to be and its total space velocity , about twice the bulk velocity of the oxygen-rich ejecta of . Their finding supports the argument of VogtDopita2011 for a polar jet. Of high significance to my claim for the JJEM is Lawetal2020 finding that the runaway knot is nearly perpendicular to the torus plane and coincident with an X-ray emission spot (inset of panel (b) of Figure 3) that is substantially enhanced in silicon and somewhat in sulfur relative to the Large Magellanic Cloud and N132D’s bulk ejecta. The different composition from the rest of the ejecta shows that the runaway knot is not a result of an instability, e.g., a finger of Rayleigh-Taylor instability that keeps the same composition stratification of the rest of the ejecta; the runaway knot must be a separate ejection event, namely, a jet. This is the same argument on why the silicon-rich jet of Cassiopeia A cannot be an instability, and it is a jet launched during the explosion process (e.g., Soker2017IAUS).
@@
Lawetal2020 find some departures from pure axi-symmetry of the torus in that the redshifted material is elevated above the midplane of the torus, while the blueshifted side is below the midplane. In the JJEM, this is part of the point-symmetric morphology that pairs of jets shape during the explosion process. Lawetal2020 also find that the runaway knot is to the normal to the torus.
@@
Lawetal2020 identified a break in the torus; I present their identification in Figure 4. Examining some images from their analysis, I identify a counter break, i.e., a break opposite to the one they identified, as shown in panel (a) of Figure 4; it has a narrower opening. In the same panel, I draw a line from the break through the center and to the counter break that I identify here. As far as I can tell, this symmetry line through the breaks in the torus coincides on the plane of the sky with the perpendicular axis that I draw in Figures 1 and 2, up to the uncertainty in the exact location of the center of explosion. I suggest that a jet, one of two jets in a pair, shaped the break in the torus. The continuation of material outward on the sides of the break supports such an interpretation. An opposite jet shaped the counter break. On panel (b) of Figure 4, I added the perpendicular axis, but shifted to the center that Lawetal2020 uses. This axis is aligned along the extension of the oxygen-rich material to the northeast, highlighting the relationship between the large-scale symmetry and the oxygen-rich material.
@@
Lawetal2020 compare the torus they reconstruct in SNR N132D with that of Cassiopeia A. I note that the torus of Cassiopeia A possesses a rich point-symmetric morphology (BearSoker2025). I here argue that the torus of SNR N132D also possesses a point-symmetric morphology. Lawetal2020 compared the runaway knot of SNR N132D to fast ejecta clumps in the Vela SNR, which is another CCSNR with a prominent and rich point-symmetric morphology that only the JJEM can account for (Soker2023SNRclass; SokerShishkin2025Vela).
The maps of N132D show some structures that are not part of a point-symmetric morphology, like the prominent bright filament extending south–southwest from the center in panels (b) and (d) of Figure 1. This is expected because several processes act to destroy point-symmetry: Instabilities at the explosion process as in the neutrino-driven simulations (see references in Section 1), one jet in a pair that is much more energetic than the opposite jet (Soker2024CounterJet; Bearetal2025Puppis), interaction with an asymmetrical CSM, the NS kick at explosion, nickel bubbles which are concentrations of nickel that when decay heat the gas and inflate bubbles, and binary star at explosion (see discussion by SokerShishkin2025Vela). This specific filament seems to result from unequal jets with instabilities that increased the asymmetry. This is for the future three-dimensional simulations of the JJEM.
@@ 4 Discussion and Summary
By visually inspecting images of CCSNR N132D, a method I justified in Section 2, I identified a point-symmetric morphology in CCSNR N132D, composed of two symmetry axes. The first one is the short symmetry axis, extending from the northwest ear and through the center of the iron-rich emission that Fosteretal2025 presented in a recent study (panel a of Figure 1). The second, the perpendicular axis on the plane of the sky, is an axis perpendicular to the short symmetry axis and passing through its center, which is also aligned along the long dimension of SNR N132D; it coincides with the extension of the oxygen-rich gas to the northeast (panel d of Figure 1, Figure 2 and 3, and panel b of Figure 4). The perpendicular axis seems to coincide with the line from the break in the torus that Lawetal2020 identify to the counter break that I identify (panel a of Figure 4). There might be a third axis: Lawetal2020 find the redshifted material in the torus to be somewhat above the midplane of the torus and the blueshifted side to be below the midplane. A line between these two regions, that is inclined to the plane of the torus, might define a third symmetry axis in SNR N132D. This deserves further study.
SNR N132D is surrounded by material (CSM and/or ISM), as indicated by the diffuse blue zones in panel (a) of Figure 3. SNR N132D interacts with the surrounding clouds (e.g., Dopitaetal2018; Sanoetal2020; HESSColl2021; Guetal2025N132D). This interaction affects the morphology of the outer regions, but cannot shape the inner oxygen-rich gas. The fact that the inner and outer morphologies correlate shows that the point symmetry is a result of the explosion.
@@
I considered the possibility that the two opposite jets along the long dimension of SNR N132D were not exactly opposite, i.e., a bent pair of jets. I mark the two suggested jets’ directions with solid-orange arrows in panel (d) of Figure 1; they are at to each other. Shishkinetal2025S147 suggested that one of the pairs of jets that exploded and shaped CCSNR S147 was a bent pair.
@@
The JJEM accounts for the point-symmetric morphology of SNR N132D, as well as for the perpendicular two axes and the bent pair of jets. The perpendicular axes might result from the following two effects. In an early energetic pair of jets, one jet carries much more momentum (and energy) than the other. As a result of that, the NS acquires a kick velocity; this is the kick-BEAP (kick by early asymmetrical pair) mechanism (Bearetal2025Puppis; Shishkinetal2025S147). The kick velocity is along the axis of the jets. The kick velocity imparts an angular momentum component to the material that the NS accretes later from the collapsing core that is perpendicular to the kick velocity. If this component dominates, then the pair of jets that the accreted gas launches along its angular momentum axis tends to be perpendicular to the kick velocity (e.g., BearSoker2018kick), i.e., perpendicular to the axis of the first pair of jets. This effect of post-kick accretion might explain the avoidance of small angles between the NS kick velocity and the main jet axis in some CCSNRs (e.g., Soker2022SNR0540; BearSoker2023RNAAS). Many of the accretion episodes through the intermittent accretion disks in the JJEM last for a time scale that is shorter, or not much longer than the relaxation time of the accretion disk (Soker2024CounterJet). Because accretion of material with stochastic angular momentum and density fluctuations forms the accretion disk, the two sides of the unrelaxed accretion disk might be unequal in size and structure. Earlier studies of the JJEM followed (Soker2024CounterJet) and considered that opposite jets differ in their power and opening angle. I here add to the claim by Shishkinetal2025S147 that, in addition, the two jets might not be exactly opposite to each other, and the angle between them can be , e.g., in SNR S147.
One of the future steps will be to conduct heavy three-dimensional hydrodynamical simulations of the expansion of the jets through the inner core, i.e., in the silicon and oxygen layers, and follow the nucleosynthesis that the jets induce. These simulations will find the kinematics of the different elements. They will account for the dynamics, e.g., the finding that the velocity dispersions of the sulfur and silicon emission lines in SNR N132D are while that of the Fe He emission line is (Xrism2024N132D)
@@
SNR N132D is the fifteenth CCSNR with an identified point-symmetric morphology attributed to the JJEM (Section 2). Presently, the point-symmetric morphology is the only property of CCSNe and CCSNRs that can clearly distinguish between the JJEM and the delayed neutrino explosion mechanism (Soker2024UnivReview). Other properties, such as light curves and nucleosynthesis, are similar but not identical between the two explosion mechanisms. However, the differences in these properties are not at a level that allows for a clear observational determination of the explosion mechanism (Soker2024UnivReview). The point-symmetrical morphology due to point-symmetric explosion is a robust prediction of the JJEM (although it will not be observed in all CCSNRs) that has no explanation in the neutrino-driven mechanism (e.g., SokerShishkin2025Vela). Some supporters of the neutrino-driven explosion mechanism have noted that their simulations do not produce jets (I attributed it to the limited ability of existing CCSN hydrodynamical numerical codes to handle magnetic field reconnection; Soker2025Learning). As a result, they often do not engage with or cite literature related to the JJEM model. This approach, where theoretical expectations (but only by some researchers) are used to dismiss observational evidence, contrasts with the standard scientific methodology in which models are tested and refined based on empirical data. In contrast, the JJEM model argues that the observed point symmetric morphologies in CCSNRs provide significant evidence that favors jet-driven processes over neutrino-driven explosions as the primary explosion mechanism. I therefore encourage the CCSNe research community to consider evaluating the JJEM on equal ground with the delayed neutrino explosion mechanism.
@@ Acknowledgements
I thank Ealeal Bear, Dima Shishkin, and Amit Kashi for their many discussions and comments, which were relevant to the topic of this study. I thank an anonymous referee for useful comments. I thank the Charles Wolfson Academic Chair at the Technion for the support.